OpenStudy (anonymous):

Verify that the given function is a solution to the differential equation: y''(x)= 5/y^3 , y = sqrt(x^2 + 5)

OpenStudy (anonymous):

I've differentiated twice and so far I've ended up with; y'' = -x^2(x^2 + 5)^(-3/2) + (x^2 + 5 )^(-1/2)

OpenStudy (anonymous):

It worked for me. I will continue from where you stopped: \[{-x^2 \over (x^2+5)^{3/2}}+{1 \over (x^2+5)^{1/2}}={-x^2+x^2+5 \over (x^2+5)^{3/2}}={5 \over ((x^2+5)^{1/2})^3}\] But \((x^2+5)^{1/2}=y\) So, \(y''=\frac{5}{y^3}\).

OpenStudy (anonymous):

\[y=\sqrt{x^2+5}=(x^2+5)^{1/2}\] \[y' = (1/2)(2x)(x^2+5)^{-1/2}=x(x^2+5)^{-1/2}\] \[y'' = (1)(x^2+5)^{-1/2} + (x)(-1/2)(2x)(x^2+5)^{-3/2}\] \[y'' = (x^2+5)^{-1/2} - x^2(x^2+5)^{-3/2}\] \[y''=(x^2+5)^1(x^2+5)^{-3/2}-x^2(x^2+5)^{-3/2}\] \[y''=x^2(x^2+5)^{-3/2}+5(x^2+5)^{-3/2}-x^2(x^2+5)^{-3/2}\] \[y''=5(x^2+5)^{-3/2}\] \[\frac{5}{y^3} = 5y^{-3} = 5((x^2+5)^{1/2})^{-3}\] \[5y^{-3}=5(x^2+5)^{-3/2}=y''\]

OpenStudy (anonymous):

ty guys :)